You bought one of Mastadon Manufacturing Co.’s 6 percent coupon bonds one year ago for $1,040. These bonds make annual payments, mature eleven years from now, and have a par value of $1,000. Suppose you decide to sell your bonds today, when the required return on the bonds is 5 percent.
  
If the inflation rate was 2.0 percent over the past year, what would be your total real return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  
Total real return             %
Answer #1

Value of the Bond today (1+r) = Coupon * - MaturityValue (1+r)

Where r is the discounting rate of a compounding period i.e. 5%

And n is the no of Compounding periods i.e. 11 years

Value of the Bond today 1- = 60 -- (1+0.05T - 0.05 - 1000 (1+0.05)11

= 498.38485314 + 584.67928905

= $ 1083.06

Nominal Return = (1083.06 - 1040) / 1040 = 4.14%

Real Rate of Return = 1+ NominalRateof Return 1+ Inflation Rate

= 1+0.0414 1+0.02

= 2.098039215% OR 2.10%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.